one billion electrons pass from wire for one microsecond then how much current will flow?
 

Grade:10

2 Answers

shashank
8 years ago
1.6x10^-4
hero
7 years ago
e = 100,00,00,000
c = 100,00,00,000*1.602*10^-19
q = 1.602*10^-10
i = q/t
t = 1 micro sec
i = 1.602*10^-10/1*10^-6
i = 1.602*10^-4